To make an international call, you need the code for the country you are calling. The codes for country A, country B, and C are three consecutive integers whose sum is 108. Find the code for each country.
Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

To make an international call, you need the code for the country you are calling. The codes for country A, country B, and C are three consecutive integers whose sum is 108. Find the code for each country.
---
1st:: x-1
2nd:: x
3rd:: x+1
------
Equation:
sum = 108
3x = 108
x = 36 (2nd)
x-1 = 35 (1st)
x+1 = 37 (3rd)
---------------
Cheers,
Stan H.
------------

Answer by addingup(3677)

A = x
B = x + 1
C = x + 2
-------------------------------
x + x + 1 + x + 2 = 108
3x + 3 = 108
3x = 105
x = 35 this is A. B = 35 + 1 = 36 and C = 35 + 2 = 37